The dimensions of aluminum foil in a box for sale in supermarket are 66 2/3 years by 12 inches. The mass of the foil is 0.83 kg. if its density is 2.70 g/cm3, then what us the thickness of the foil in inches?


1
Expert's answer
2021-02-02T01:26:16-0500

First convert the all dimesions in required

66 2/3yard*(36inch/1 yard) = 2400 inches

0.83kg= 830gm

volume of the foil= l*b*h

volume= 2400*12*h

density=2.7g/cm"^3" = 44,245g/inch"^3"

"density= \tfrac{mass}{volume}"

substitute all values

"44,245= \tfrac{830}{2400*12*h}"

h= 0.00065inch